OpenStudy (anonymous):

using the point-slope form y-y1=m(x-x1) how do i write a equation for a line with a slope of -2, containing the point (3,4)

OpenStudy (anonymous):

(3,4) is (x1,y1), and -2 is m, so just plug that in y-4=-2(x-3) y=-2x+6+4 y=-2x+10